A certain gas is present in a 15.0 L cylinder at 2.0 atm pressure. If the pressure is increased to 4.0 atm , the volume of the gas decreases to 7.5 L . Find the two constants ki, the initial value of k, and kf, the final value of k, to verify whether the gas obeys Boyle’s law.
Answer #1

Boyle's Law states that at constant temperature volume of a gas is inversely proportional to pressure.Or mathematically

PV = k (constant).

For 1st case PV = 15*2 =30 = ki

For 2nd case PV= 7.5*4 = 30 = kf

So both the constants are same . So boyes law is verified.